OpenStudy (anonymous):

he ratio of two numbers is 5 to 3. The sum of the numbers is 16. What are the numbers?

OpenStudy (anonymous):

answers could be 8 and 8 9 and 7 15 and 9 10 and 6

OpenStudy (anonymous):

@Easyaspi314 help please.

hartnn (hartnn):

let the 2 numbers be x and y so, ratio = x/y = 5/3 got this ? sum = 16 can you make other equation ?

OpenStudy (anonymous):

so the answer would be 9 and 7

hartnn (hartnn):

no...the ratio for 9 to 7 is not 5 to 3 ...

hartnn (hartnn):

you were trying by using choices ? tried last choice ?

OpenStudy (anonymous):

ohhhh i get it its 10 and 6

hartnn (hartnn):

yes! :)

OpenStudy (anonymous):

thanks for your help! @hartnn

hartnn (hartnn):

you're welcome ^_^